please help me solve this equation.
3x+y=-3
2x+y=-1

Answer by richwmiller(17219)

There are two equations. It is called a system of equations.
Solve both for y and set them equal
y=-3-3x
y=-1-2x
-3-3x=-1-2x
add 3x to both sides
-3=-1+x
add 1 to both sides
-2=x
2(-2)+y=-1
-4+y=-1
y=3
(-2,3)